brian asked a group of people their favourite holiday destination.

the results are summarised in the table.


destination uk. europe. usa. africa. other

frequency. 156. 270. 204. 216. 234

angle. ? ? ? ? ?


work out the size of each angle to draw a pie chart


a= ? b= ? c=? d= ? e= ?


1
Expert's answer
2021-05-10T13:03:05-0400

Total number of people:

"156+270+204+216+ 234=1080"


So, angle in pie-chart will be:


UK:


"\frac{156}{1080}\cdot360\degree=52\degree"



Europe:


"\frac{270}{1080}\cdot360\degree=90\degree"



USA:


"\frac{204}{1080}\cdot360\degree=68\degree"



Africa:


"\frac{216}{1080}\cdot360\degree=72\degree"



Other:


"\frac{234}{1080}\cdot360\degree=78\degree\\\"

Pie-Chart: